# TESSERA: Temporal Embeddings of Surface Spectra for Earth Representation and Analysis

### v1

@@ 2.3 Canopy height estimation
Canopy height is a key structural attribute of forests, closely linked to above-ground biomass and carbon stocks, and therefore important for climate mitigation efforts and carbon accounting frameworks like REDD+ [57, 26]. It reflects ecosystem function [56], forest age, and successional stage, while also serving as a proxy for habitat quality and vertical complexity—important determinants of biodiversity patterns and species richness [20, 15, 73]. Accurate, scalable canopy height models (CHMs) enable detection of forest degradation and disturbance arising from selective logging, disease outbreaks and storm damage.
@@
Despite substantial progress in global canopy height mapping, current approaches exhibit limitations that restrict their reliability and applicability. GEDI, a spaceborne LiDAR system, provides spatially sparse measurements of canopy height that are used to train models for generating maps from optical and radar data; however, GEDI does not sample high latitudes [21], and produces unreliable estimates in mountainous terrain [27]. Radar-based CHMs (e.g., TanDEM-X) achieve global coverage but are affected by canopy height saturation in dense forests and depend on high-quality terrain models to isolate canopy height accurately. Optical data, while widely available, lacks inherent sensitivity to vertical structure and requires indirect proxies or fusion with information from active sensors  [82, 50] - yet none of the three most widely used global CHM models currently implement sensor fusion [46, 68, 78]. Most approaches instead rely on a single sensor type and derive height predictions either from temporal composites or from large numbers of hand-engineered features (e.g., $>500$ features in Potapov et al. [68]), which may generalize poorly across ecosystems. These models tend to under-predict tall canopies, especially in tropical forests where the majority of above-ground carbon is stored, and face challenges from persistent cloud cover, sensor saturation [45], and inconsistencies in input quality. As a result, significant discrepancies remain between global CHMs and airborne LiDAR benchmarks, with current products underestimating local variability and structural extremes [58]. Most CHMs are also static, failing to capture seasonal or interannual dynamics, and inconsistencies in methodology and sensor characteristics hinder cross-product integration. These limitations underscore the need for scalable, temporally resolved, and sensor-fused CHMs with quantified uncertainties to support forest monitoring, carbon accounting, and Earth system modelling [59, 65].
@@
We evaluated TESSERA representations to predict the height of the airborne LiDAR-derived canopy at 10 m resolution within a 5 × 6 km area of tall old-growth tropical forest in the Danum Valley, Borneo[40]. To ensure spatial independence, we conducted four-fold spatial cross-validation, each time holding out a contiguous 50% of the region for testing and using the remaining 50% for training and validation. A 30-million-parameter UNet was trained using 64 × 64 pixel patches of the representations as input. To account for variability in training convergence, we performed three independent training runs per fold, each for 200 epochs, and retained the model with the best performance on the validation set. Accuracy metrics for the foundation models are reported as the average across all 12 runs. TESSERA achieved an $R^{2}$ of 0.66, a root mean squared error (RMSE) of 8.88 m, and a mean bias of $-0.62$ m.
@@
We benchmarked TESSERA’s performance against state-of-the-art approaches through three comparisons (Fig. 3). First, we compared predictions against three global canopy height products: Potapov et al. [68] (‘GLAD’ map), Lang et al. [46] (‘ETH’ map), and Tolan et al. [78] (‘Meta’ map). Although local models generally outperform global products in regional evaluations [70], the pronounced saturation of these global models at much lower canopy heights—combined with RMSEs exceeding their reported global averages [68, 46, 78]—underscores the challenge of accurately mapping structurally complex, tall-canopy regions such as Danum Valley. Second, we compared TESSERA to the PRESTO [79] foundation model, using the same model architecture and training procedure but replacing the inputs of TESSERA with the representations of PRESTO. Third, we compared our results to regional wall-to-wall canopy height maps for Indonesia, Malaysia, and the Philippines generated by Lang et al. [47], which used deep convolutional neural networks trained on Sentinel-2 imagery and GEDI LiDAR reference data for regional canopy height and biomass estimation.
@@
In all comparisons, TESSERA outperformed competing methods. Moreover, its representations do not require preprocessing, greatly enhancing reproducibility. Instead, only the region and year of interest are needed to replicate the input data pipeline. By fusing Sentinel-1 and Sentinel-2 data without requiring preprocessing or feature engineering, TESSERA’s 128-dimensional representations encode cloud-free, information-rich signals that effectively address key challenges in canopy height modelling.
Figure 3: TESSERA representations outperform global and regional models in canopy height estimation. Top row: Density scatter plots compare predicted canopy heights from the TESSERA (left) and PRESTO (middle) models against airborne LiDAR-derived heights. TESSERA achieves higher accuracy than PRESTO, demonstrating superior performance in capturing structural forest attributes. The rightmost panel shows model bias across 10m canopy height bins, with LiDAR values binned on the x-axis and model predictions on the y-axis. TESSERA more closely follows the 1:1 line, particularly in mid- to high-canopy regimes, while PRESTO underestimates heights above  40m. Histogram above indicates pixel count distribution per bin. Bottom row: Comparison with three global canopy height products— ETH Global CHM, GLAD (Potapov) Global CHM, Meta (Tolan) Global CHM, and ETH Regional CHM—evaluated on the same test region. All four products exhibit substantially lower predictive performance ($R^{2}<0.05$, $RMSE>14m$), including negative correlations for three out of four models. These results underscore the difficulty of accurately estimating tall tropical canopies using global models trained without local data. Rightmost column: Spatial comparison between true canopy height (LiDAR) and TESSERA-predicted canopy height for a representative test patch, illustrating the model’s capacity to capture fine-scale structural variation while still saturating for taller canopy values.
@@ 2.4 Burned Area Detection
Wildfire frequency and severity are projected to increase in many parts of the world due to climate change [3, 39, 66]. Accurate maps of burned areas (BAs) and burn severity are critical for monitoring wildfire trends [6], predicting wildfire occurrence [17], assessing ecosystem recovery [89], and establishing effective fire management strategies [33]. Remote sensing is widely used to detect burned areas and assess severity regionally and globally [44].
@@
Most existing BA mapping approaches rely on time series analysis or direct comparison between pre-fire and post-fire images, often requiring careful selection of suitable cloud-free images, manually tuned thresholds, and expert input [44, 67]. These methods typically use spectral indices such as the Normalized Burn Ratio (NBR), though many studies have also explored the design of custom features and applied spectral dimensionality reduction techniques to preserve key fire-related signals [44]. The majority of existing methods use optical sensors [44], which can detect fire-specific spectral signatures but are susceptible to cloud cover [34], particularly problematic for fire detection in ecosystems where vegetation rapidly re-greens [35]. SAR offers complementary sensitivity to vegetation structure and is unaffected by cloud cover [44, 76, 81]. Recent reviews emphasize the need for sensor fusion, combining passive and active sensors to improve accuracy and robustness [44, 18]. Sentinel sensors in particular, when combined together at 10 m resolution, have shown promise in detecting BA from small fires that are often left out of coarse, global BA products [76, 81, 71]. TESSERA extends the philosophy of traditional approaches by encoding temporal dynamics, while addressing major limitations through sensor fusion and the use of abstract, data-driven features learned directly from large volumes of combined Sentinel-1 and Sentinel-2 imagery (Fig. 4).
@@
Figure 4: TESSERA representations capture wildfire extent, timing, and severity across two burned regions. (a–c) Burn Area A, which experienced two distinct fires in June and August 2021. (a) Time of fire map based on MTBS, showing unburned areas (blue), the June fire (orange), and the August fire (red). (b) Burn severity classification for the same region, with unburned (blue), low (green), moderate (orange), and high (red) severity. (c) UMAP projection of TESSERA embeddings from 2021, with colors corresponding to burn severity. Embeddings form well-separated clusters for unburned, June-, and August-burned areas, with a visible internal gradient reflecting severity within August fire. (d–f) Burn Area B, which experienced a single large fire in September 2021. (d) Time of fire map showing unburned (blue) and burned (green) areas. (e) Burn severity classification showing heterogeneous severity across the landscape. (f) UMAP projection for Burn Area B, revealing fuzzier boundaries between burned and unburned pixels, though the severity gradient remains evident.
@@
We tested whether TESSERA’s temporal representations capture fire-induced changes in spectral and radar signals. We evaluated its ability to detect burned areas and burn severity using two randomly selected 15 × 15 km regions from the Burn Scar Benchmarking dataset, each affected by wildfires in California during 2021. TESSERA representations for 2021 were obtained and UMAP dimensionality reduction was applied to a random subsample of pixels. Fire boundaries and burn severity data were sourced from the Monitoring Trends in Burn Severity (MTBS) dataset[2]. The resulting UMAP projections demonstrated strong separability in two-dimensional space, visually highlighting TESSERA’s capacity to detect disturbances at the pixel level. Burn Area A experienced two distinct wildfires—in June and August—while Burn Area B underwent a single, large fire affecting mostly forested areas, resulting in heterogeneous burn severity. The UMAP revealed separability across three domains (Fig. 4):
@@
  • Burned area detection: In both regions, burned pixels formed distinct clusters from unburned ones, with clearer separation in Burn Area A.
@@
  • Burn severity gradient: In both regions, a gradient structure in UMAP space corresponded to burn severity, indicating that TESSERA encodes disturbance magnitude.
Because TESSERA encodes annual time series of spectral and radar signals for each pixel, it inherently captures changes in pixel trajectories without requiring prior preprocessing or the selection of cloud-free before-and-after images. This capacity would support a simplified, scalable pipeline for large-area disturbance monitoring.
2.5 Above-ground biomass estimation
Forests play a central role in the global carbon cycle, creating a pressing need for reliable and scalable AGB estimates for climate modeling, carbon accounting, and informed strategies for conservation and sustainable land use [36]. Accurate estimation of above-ground biomass (AGB) is essential for quantifying forest carbon stocks and monitoring temporal changes due to deforestation, degradation, and recovery [22].
We evaluated the performance of TESSERA in predicting AGB against two leading models on a benchmark dataset. Firstly, we considered the geospatial foundation model SpectralGPT [37], which was found in the Pangaea benchmark [53] to outperform other tested foundation models on the BioMassters dataset [60]. Secondly, as the state-of-the-art bespoke model, we considered the winner of the Biomassters competition, a UNet with a temporal attention encoder [60]. The BioMassters dataset comprises 11,462 image patches representing above-ground biomass across Finnish forests from 2017 to 2021. Ground-truth AGB values are derived from airborne LiDAR and aerial imagery, using calibrated allometric equations informed by extensive field plots by the Finnish Forest Centre and National Land Survey. Each patch covers a 2560×2560 m area, with 10×10 m spatial resolution per pixel. The fine spatial granularity of the dataset makes the estimation task more challenging, as it introduces a long, thin tail of high AGB values that would not appear in coarser-resolution data.
A comparison of RMSE across varying training label fractions (Fig. 5a) shows that a UNet trained on the TESSERA representations consistently surpasses a UPerNet trained on the SpectralGPT representations, achieving lower prediction errors on 1% of the training set than SpectralGPT on the full training set. These results underscore the robustness of TESSERA in low-label regimes. Moreover, while the best task-specific model trained on the full dataset outperforms TESSERA, the gap is moderate, even with low label availability. In general, bespoke models are inherently difficult to surpass in fully supervised settings. It is worth noting that the comparison to benchmarks is not entirely straightforward, as the satellite data used in the benchmark models differs from the TESSERA input. However, these discrepancies can be considered as an inherent part of the respective processing pipelines and hence largely unavoidable.
Figure 5: TESSERA representations demonstrate robustness under limited label availability. a, TESSERA achieves lower RMSE values than the foundation model benchmark SpectralGBT (RMSE values from Ref. [53]). Following the evaluation protocol of Ref. [53], RMSE is computed on a fixed test set of 2773 patches, with the remaining data randomly split into training and validation sets (80:20). The UNet model is trained on the TESSERA representations for 80 epochs, with the checkpoint achieving the lowest validation loss selected as the final model. For limited labels testing, a subset of labels is randomly chosen from the training set. The inset depicts the coefficient of determination ($R^{2}$) and mean bias (MB) as functions of label availability. b, Predicted AGB map for an example patch using 10% of the training labels. c, The mean predicted AGB values show increasing bias at higher ground-truth AGB values, which are sparse in the dataset. d, Pixel-wise comparison of predicted versus ground-truth AGB values for the test data with TESSERA+UNet (top) and Biomassters winner (bottom).

### v2

@@ 2.2 Canopy height estimation
Canopy height is a key structural attribute of forests, linked to above-ground biomass and carbon stocks, and therefore important for climate mitigation efforts and carbon accounting frameworks such as REDD+ [106, 49]. It is correlated with ecosystem function [105], forest age, and successional stage, while also serving as a proxy for habitat quality, an important determinant of biodiversity patterns and species richness [37, 25, 136]. Accurate, scalable canopy height models (CHMs) can enable detection of forest degradation and disturbance.
@@
Despite substantial progress in global canopy height mapping, current approaches exhibit limitations that restrict their reliability and applicability. GEDI, a spaceborne LiDAR system, can provide reasonably accurate canopy height measurements to train models to generate CHMs from optical and radar data; however, it does not provide wall-to-wall maps, and does not sample high latitudes [41], and produces unreliable estimates in mountainous terrain [52]. Radar-based CHMs (e.g., TanDEM-X) achieve global coverage, but are affected by canopy height saturation in dense forests and depend on high-quality terrain models to accurately isolate canopy height [131]. Optical data, while widely available, lack inherent sensitivity to vertical structure and require indirect proxies or fusion with information from active sensors  [158, 96] - yet none of the three most widely used global CHM models currently implement sensor fusion [84, 121, 148]. Most approaches instead rely on a single sensor type and derive height predictions either from temporal composites or from large numbers of hand-engineered features (e.g., >500>500> 500 features in Potapov et al. [121]), which may generalize poorly across ecosystems. These models tend to underestimate tall canopies, especially in tropical forests where most of the above-ground carbon is stored, and face challenges from persistent cloud cover, sensor saturation [83], and inconsistencies in input quality. As a result, significant discrepancies remain between global CHMs and airborne LiDAR benchmarks, and current products underestimate local variability and structural extremes [107]. Most CHMs are also static, failing to capture inter-annual dynamics, and inconsistencies in methodology and sensor characteristics hinder cross-product integration. These limitations underscore the need for scalable, temporally resolved, and sensor-fused CHMs with quantified uncertainties to support forest monitoring, carbon accounting, and Earth system modelling [108, 116].
@@
We evaluated TESSERA embeddings to predict canopy height at 10 m resolution within a 5 $\times$ 6 km area of tall old-growth tropical forests in the Danum Valley, Borneo [74] with openly available ground-truth data [33]. To ensure spatial independence, we conducted four-fold spatial cross-validation against the airborne LiDAR-derived canopy height, each time holding out a contiguous 50% of the region for testing and using the remaining 50% for training and validation. A 30-million-parameter UNet was trained using 32 $\times$ 32 pixel patches of the embeddings as input. To account for variability in training convergence, we performed three independent training runs per fold, each for 200 epochs, and retained the model with the best performance on the validation set. The accuracy metrics for the foundation models are reported as the average across the 12 runs. TESSERA achieved an $R^{2}$ of 0.66, a root mean squared error (RMSE) of 8.88 m, and a mean bias of 0.62-0.62- 0.62 m.
@@
We benchmarked TESSERA’s performance against state-of-the-art approaches through three sets of comparisons (Fig. 3). First, we evaluated its predictions against three global canopy height products: the GLAD map by Potapov et al. [121], the ETH map by Lang et al. [84], and the Meta map by Tolan et al. [148]. Although local models generally outperform global products in regional evaluations [127], the pronounced saturation of these global models at much lower canopy heights - combined with RMSEs that exceed their reported global averages [121, 84, 148]—underscores the challenge of accurately mapping structurally complex high canopy regions such as Danum Valley.
We also compared our results to regional wall-to-wall canopy height maps for Indonesia, Malaysia, and the Philippines produced by Lang et al. [85], which used deep convolutional neural networks trained on Sentinel-2 time series and GEDI LiDAR to estimate regional canopy height and biomass. Second, we compared TESSERA to two alternative pixel-based foundation models, PRESTO [149] and Google Satellite Embedding (GSE) [56], using the same U-Net architecture, but replacing TESSERA’s input representations with those of the other foundation models. Finally, to assess the added value of the learned representations over the raw input data, we trained the same U-Net using cloud-free annual median composites of Sentinel-2 imagery.
@@
In all comparisons, TESSERA outperformed competing methods. Moreover, its embeddings do not require pre-processing, greatly enhancing reproducibility. Instead, only the region and year of interest are needed to replicate the input data pipeline. By fusing Sentinel-1 and Sentinel-2 data without requiring preprocessing or feature engineering, TESSERA’s 128-dimensional embeddings encode cloud-free, information-rich signals that effectively address key challenges in canopy height modelling.
Figure 3: TESSERA outperforms global and regional models as well as alternative pixel-based foundation models in canopy height estimation. Predicted canopy height (y-axis) is compared against airborne LiDAR reference data (x-axis) across multiple models and input datasets in Danum Valley, Borneo. Top row: Predictions from models trained using different input representations—TESSERA, PRESTO, Google Satellite Embedding (GSE), and raw Sentinel-2 composites—using the same model architecture and training procedure. TESSERA achieves the highest coefficient of determination and the lowest RMSE (8.88±0.988.88$\pm$ 0.988.88 ± 0.98 m), indicating superior performance in structurally complex, tall tropical canopies. Bottom row: Comparisons to existing global and regional canopy height maps, including products from ETH, GLAD, Meta, and Lang et al.’s regional CHM. All four products show substantially lower predictive performance ($R^{2}<0.05$, $RMSE>14$ m), with three exhibiting negative correlations. These models saturate at lower canopy heights and show large bias and RMSE, highlighting the limitations of global products when applied to high-biomass tropical forests without local calibration. A 1:1 reference line (blue) is shown in each panel for comparison.
2.3 Above-ground biomass estimation
Forests play a central role in the global carbon cycle, creating a pressing need for reliable and scalable AGB estimates for climate modeling, carbon accounting, and informed strategies for conservation and sustainable land use [66]. Accurate estimation of above-ground biomass (AGB) is essential to quantify forest carbon stocks and monitor temporal changes due to deforestation, degradation, and recovery [45].
We evaluated the performance of TESSERA in predicting AGB against two leading models on a benchmark dataset. Firstly, we considered the geospatial foundation model SpectralGPT [67] combined with a UPerNet regression head – a setup found in the Pangaea benchmark [101] to outperform other foundation models tested on the BioMassters dataset [111]. As the preprocessing pipeline is not yet available in the codebase of Ref. [67] at the time of our analysis, we adopted the implementation from Ref. [101], which uses the preprocessed satellite imagery from the BioMassters dataset as input. Secondly, we used the winning model from the BioMassters competition as our state-of-the-art supervised baseline – a UNet with a temporal attention encoder [111], which achieved the lowest RMSE among over 1,000 submissions. The BioMassters dataset comprises 11,462 image patches that represent above-ground biomass in Finnish forests from 2017 to 2021. Ground-truth AGB values are derived from airborne LiDAR and aerial imagery, using calibrated allometric equations informed by extensive field plots by the Finnish Forest Centre and National Land Survey. Each patch covers a $2560\times 2560\,\textrm{m}^{2}$ area, with a spatial resolution of 10 m per pixel. The fine spatial granularity of the dataset makes the estimation task more challenging, as it introduces a long, thin tail of high AGB values that would not appear in coarser-resolution data.
A comparison of RMSE across varying training label fractions (Fig. 4a) shows that a UNet trained on the TESSERA embeddings consistently surpasses a UPerNet trained on the SpectralGPT representations, achieving lower prediction errors on 1% of the training set than SpectralGPT on the full training set. These results underscore the robustness of TESSERA in sparse-label regime. Moreover, TESSERA achieves results comparable to those of the supervised baseline trained on the full dataset, even with limited label availability. In general, bespoke models are inherently difficult to surpass in fully supervised settings. It is worth noting that the comparison to benchmarks is not entirely straightforward, as the satellite data used in the benchmark models differs from the TESSERA input. However, these discrepancies can be considered as an inherent part of the respective processing pipelines and hence largely unavoidable.
Figure 4: TESSERA embeddings demonstrate robustness for AGB mapping under limited label availability. a, TESSERA achieves lower RMSE values than the foundation model benchmark SpectralGPT. Following the evaluation protocol in Ref. [111], RMSE is computed on a fixed test set of 2773 patches and reported as the average RMSE per patch. For limited labels testing, a subset of labels is randomly chosen from the remaining data, and randomly split into training and validation sets (80:20). The figure shows the average RMSE computed over five random splits, with bars indicating the minimum–maximum range when it exceeds the size of the dot. The UNet model is trained on the TESSERA embeddings for 80 epochs, with the model achieving the lowest loss on the validation set selected as the final model. The inset depicts the coefficient of determination ($R^{2}$) and mean bias (MB) as functions of label availability. b, Predicted AGB map for an example patch using 10% of the training labels. c, The mean predicted AGB values show increasing bias at higher ground-truth AGB values, which are sparse in the dataset. The top panel depicts the histogram of the ground-truth AGB values. d, Pixel-wise comparison of predicted versus ground-truth AGB values for the test data with TESSERA+UNet (top), the supervised baseline (middle) and SpectralGPT+UPerNet (bottom).
@@ 2.4 Burned Area Detection
The frequency and severity of wildfires are projected to increase in many parts of the world due to climate change [5, 71, 117]. Accurate maps of burned areas and burn severity are critical for monitoring wildfire trends [10], predicting wildfire occurrence [29], assessing ecosystem recovery [186], and establishing effective fire management strategies [63]. Remote sensing is widely used to detect burned areas and assess severity regionally and globally [82].
@@
Most existing burn area mapping approaches rely on time series analysis or direct comparison between pre-fire and post-fire images, often requiring careful selection of suitable cloud-free images, manually tuned thresholds, and expert input [82, 119]. These methods typically use spectral indices such as the Normalized Burn Ratio (NBR), though many studies have also explored the design of custom features and applied spectral dimensionality reduction techniques to preserve key fire-related signals [82]. Most existing methods use optical sensors [82], which can detect fire-specific spectral signatures but are susceptible to cloud cover [64], particularly problematic for fire detection in ecosystems where vegetation rapidly re-greens [65]. SAR offers complementary sensitivity to vegetation structure and is unaffected by cloud cover [82, 143, 154]. Recent reviews emphasize the need for sensor fusion, combining passive and active sensors to improve accuracy and robustness [82, 30]. Sentinel-1 and 2 sensors in particular, when combined together at 10 m resolution, have shown promise in detecting burn areas from small fires that are often left out of coarse, global burn area products [143, 154, 129]. TESSERA extends the philosophy of traditional approaches by encoding temporal dynamics, while addressing major limitations through sensor fusion and the use of data-driven features learned directly from large volumes of combined Sentinel-1 and Sentinel-2 imagery (Fig. 5).
@@
Figure 5: TESSERA embeddings capture wildfire extent, timing, and severity across two burned regions. (a–c) Burn Area A, which experienced two distinct fires in June and August 2021. (a) Time of fire map based on MTBS, showing unburned areas (blue), the June fire (orange), and the August fire (red). (b) Burn severity classification for the same region, with unburned (blue), low (green), moderate (orange), and high (red) severity. (c) UMAP projection of TESSERA embeddings from 2021, with colors corresponding to burn severity. Embeddings form well-separated clusters for unburned, June-, and August-burned areas, with a visible internal gradient reflecting severity within August fire. (d–f) Burn Area B, which experienced a single large fire in September 2021. (d) Time of fire map showing unburned (blue) and burned (green) areas. (e) Burn severity classification showing heterogeneous severity across the landscape. (f) UMAP projection for Burn Area B, revealing fuzzier boundaries between burned and unburned pixels, though the severity gradient remains evident.
@@
We tested whether TESSERA’s temporal embeddings capture fire-induced changes in spectral and radar signals. We evaluated its ability to detect burned areas and burn severity using two randomly selected 15 $\times$ 15 km regions from the Monitoring Trends in Burn Severity (MTBS) dataset [4], each affected by wildfires in California during 2021. TESSERA embeddings for 2021 were obtained and UMAP dimensionality reduction was applied to a random subsample of pixels. Fire boundaries and burn severity data were sourced from the dataset. The resulting UMAP projections demonstrated strong separability in two-dimensional space, visually highlighting TESSERA’s capacity to detect disturbances at the pixel level. Burn Area A experienced two distinct wildfires—in June and August—while Burn Area B underwent a single large fire that affected mostly forested areas, resulting in a heterogeneous severity of the burn. The UMAP revealed separability across three domains (Fig. 5):
@@
  • Burned area detection: In both regions, burned pixels formed distinct clusters from unburned, with a clearer separation in Burn Area A.
@@
  • Burn severity gradient: In both regions, a gradient structure in the UMAP space corresponded to burn severity, indicating that TESSERA encodes the magnitude of the disturbance.
Because TESSERA encodes an annual time series of spectral and radar signals for each pixel, it inherently captures changes in pixel trajectories without requiring prior preprocessing or the selection of cloud-free before-and-after images. This capacity supports a simplified, scalable pipeline for monitoring large-area disturbances.